Fortune Cookie Inc. (FCI) issued $10 million of 10-year, 5% convertible bonds on April 1, Year 5 at 102.5 Coupons are payable on April 1 and October 1. Bonds without conversion privileges would have sold at 101.5. FCI’s fiscal year-end is December 31. Please assume that FCI follows IFRS.

(1) On October 1, Year 10, 30% of these bonds were converted to common shares right after the payment of interest. Determine the amount to be assigned to common shares at the time of conversion. [7 marks]

(2) On December 31, Year 10, additional 30% of these bonds were converted to common shares. Accrued interest was paid at the time of conversion. Determine the amount to be assigned to common shares at the time of conversion.
Answer #1

1. In case FCI follows IFRS the amount to be assigned to common shares at the time of conversion will be:

• 3 million bonds were converted to 30,000 common shares

• 3 million bonds were sold at102 per bond

•Thus, total proceeds from 3 million bonds

=102×3= 306 million

•present value of the Cash inflows will be:

=$ 306 million x 0.711 (5% for 7 years)

= $ 217.566 million

2. Therefore following IFRS, the 30,000 Common Shares will be valued at the fair value of the Convertable debt instrument i.e. the bonds -the present value of the Cash in-flows = 306-217.566 million= $88.434 Million.